# Differentiable Pareto-Smoothed Weighting for High-Dimensional Heterogeneous Treatment Effect Estimation

### v2

@@ Abstract
There is a growing interest in estimating heterogeneous treatment effects across individuals using their high-dimensional feature attributes. Achieving high performance in such high-dimensional heterogeneous treatment effect estimation is challenging because in this setup, it is usual that some features induce sample selection bias while others do not but are predictive of potential outcomes. To avoid losing such predictive feature information, existing methods learn separate feature representations using inverse probability weighting (IPW). However, due to their numerically unstable IPW weights, these methods suffer from estimation bias under a finite sample setup. To develop a numerically robust estimator by weighted representation learning, we propose a differentiable Pareto-smoothed weighting framework that replaces extreme weight values in an end-to-end fashion. Our experimental results show that by effectively correcting the weight values, our proposed method outperforms the existing ones, including traditional weighting schemes.
@@ 1 Introduction
To fulfill such important but often overlooked needs, several data-driven feature separation methods have been proposed [Kuang et al., 2017, Hassanpour and Greiner, 2020, Kuang et al., 2020, Wang et al., 2023]. Among them, disentangled representations for counterfactual regression (DRCFR) [Hassanpour and Greiner, 2020] aims to avoid losing useful feature information for heterogeneous treatment effect estimation by learning different representations for confounders and adjustment variables. To achieve this, this method employs a technique of inverse probability weighting (IPW) [Rosenbaum and Rubin, 1983], which performs weighting based on the inverse of the probability called a propensity score. However, such weights often take extreme values (especially in high-dimensional setups [Li and Fu, 2017]), and even alight propensity score estimation error may lead to large treatment effect estimation error.
To resolve this issue, we develop a weight correction framework that utilizes a technique in extreme value statistics, called Pareto smoothing [vehtari2015pareto], which replaces the extreme weight values with the quantiles of generalized Pareto distribution (GPD). Indeed, Zhu et al. [2020] already adopted Pareto smoothing and empirically showed that it can construct a numerically stable estimator of the average treatment effect (ATE) over all individuals.
@@ 2.1 Problem Setup
Our goal is to achieve a high CATE estimation performance in a high-dimensional setup, where the number of features $d$ is relatively large. Under this setup, removing the sample selection bias by transforming all features X into a single balanced representation might be overly severe, which leads to inaccurate treatment effect estimation. Although several representation learning methods aim to avoid such an overly severe balancing [Kuang et al., 2017, Wang et al., 2023], most are designed for ATE, not for CATE. By contrast, the DRCFR method [Hassanpour and Greiner, 2020] focuses on CATE estimation and effectively balances high-dimensional features X by weighted representation learning.
@@ 3.1 Overview
To resolve this issue, we improve the weight stability by replacing an extreme value of $w_{i}$ in Eq. (4). with a weight stabilization technique, called Pareto smoothing.
@@ 3.3 Non-Differentiable Procedures
One may consider a separate learning approach that trains the propensity score model, computes the Pareto-smoothed IPW weights by (21), and learns the feature representations. This approach, however, requires directly fitting a propensity score model to features X, not their representations. Since accurately estimating a propensity score model from high-dimensional features X is considerably difficult, due to the model misspecification error, such a separate learning approach leads to CATE estimation bias. We experimentally show its poor performance in Section 4.1.
@@ 3.4.2 Reformulation of GPD Parameter Estimators
To employ differentiable rank $\textbf{{r}}=r_{\varepsilon}(\textbf{{w}})$ for Pareto smoothing, since it takes continuous values, we need to modify the GPD parameter estimators, i.e., $\hat{\mu}$, $\hat{\sigma}$, and $\hat{\xi}$.
@@ 3.4.3 Overall Algorithm
Compared with DRCFR [Hassanpour and Greiner, 2020], our method requires additional time to compute Pareto-smoothed weights (lines 12-16 in Algorithm 1). In particular, computing differentiable rank (line 12) requires time complexity $\mathcal{O}(B\mathrm{log}B)$ for mini-batch size $B$, which is needed to evaluate the objective function in (3) and its gradient for each iteration in the training phase.
Remark: Strictly speaking, the choice of activation functions in propensity score $\pi$ and feature representations $\Gamma$ and $\Delta$ is critical for satisfying the assumption of Pareto smoothing that the distribution of the importance sampling weight is absolutely continuous, which is employed to prove the asymptotic consistency in Theorem 1 of Vehtari et al. [2024]. This assumption holds if each activation is differentiable almost everywhere (i.e., differentiable except on a set of measure zero). However, for instance, using the rectified linear unit (ReLU) in propensity score model $\pi$ makes the distribution of IPW weight discontinuous, thus violating the assumption of Pareto smoothing. Even with almost everywhere differentiable activation functions, due to the lack of learning theory on neural network models, deriving the asymptotic consistency of our CATE estimator is extremely challenging and is left as our future work.
@@ B.1 Settings of Baselines
We evaluated the performance of the neural network methods (TARNet and GANITE), with the existing implementations [Curth and van der Schaar, 2021, Yoon et al., 2018].  [^4] https://github.com/AliciaCurth/CATENets  [^5] https://github.com/jsyoon0823/GANITE By contrast, we implemented the DRCFR method using PyTorch [Paszke et al., 2019].

### v4

@@ Abstract
There is a growing interest in estimating heterogeneous treatment effects across individuals using their high-dimensional feature attributes. Achieving high performance in such high-dimensional heterogeneous treatment effect estimation is challenging because in this setup, it is usual that some features induce sample selection bias while others do not but are predictive of potential outcomes. To avoid losing such predictive feature information, existing methods learn separate feature representations using inverse probability weighting (IPW). However, due to their numerically unstable IPW weights, these methods suffer from estimation bias under a finite sample setup. To develop a numerically robust estimator by weighted representation learning, we propose a differentiable Pareto-smoothed weighting framework that replaces extreme weight values in an end-to-end fashion. Our experimental results show that by effectively correcting the weight values, our proposed method outperforms the existing ones, including traditional weighting schemes. Our code is available at https://github.com/ychika/DPSW.
@@ 1 Introduction
To fulfill such important but often overlooked needs, several data-driven feature separation methods have been proposed [Kuang et al., 2017, Hassanpour and Greiner, 2020, Kuang et al., 2020, Wang et al., 2023]. Among them, disentangled representations for counterfactual regression (DRCFR) [Hassanpour and Greiner, 2020] aims to avoid losing useful feature information for heterogeneous treatment effect estimation by learning different representations for confounders and adjustment variables. To achieve this, this method employs a technique of inverse probability weighting (IPW) [Rosenbaum and Rubin, 1983], which performs weighting based on the inverse of the probability called a propensity score. However, such weights often take extreme values (especially in high-dimensional setups [Li and Fu, 2017]), and even slight propensity score estimation error may lead to large treatment effect estimation error.
To resolve this issue, we develop a weight correction framework that utilizes a technique in extreme value statistics, called Pareto smoothing [Vehtari et al., 2024], which replaces the extreme weight values with the quantiles of generalized Pareto distribution (GPD). Indeed, Zhu et al. [2020] already adopted Pareto smoothing and empirically showed that it can construct a numerically stable estimator of the average treatment effect (ATE) over all individuals.
@@ 2.1 Problem Setup
Our goal is to achieve a high CATE estimation performance in a high-dimensional setup, where the number of features $d$ is relatively large. Under this setup, removing the sample selection bias by transforming all features X into a single balanced representation might be overly severe, which leads to inaccurate treatment effect estimation. Although several representation learning methods aim to avoid such an overly severe balancing [Kuang et al., 2017, Wang et al., 2023], most are designed to estimate ATE, not CATE. By contrast, the DRCFR method [Hassanpour and Greiner, 2020] focuses on CATE estimation and effectively balances high-dimensional features X by weighted representation learning.
@@ 3.1 Overview
To resolve this issue, we improve the weight stability by replacing an extreme value of $w_{i}$ in Eq. (4) with a weight stabilization technique, called Pareto smoothing.
@@ 3.3 Non-Differentiable Procedures
One may consider a separate learning approach that trains the propensity score model, computes the Pareto-smoothed IPW weights by (21), and learns the feature representations. This approach, however, requires directly fitting a propensity score model to features X, not their representations. Since accurately estimating a propensity score model from high-dimensional features X is considerably difficult, such a separate learning approach yields large model misspecification error and hence leads to CATE estimation bias. We experimentally show its poor performance in Section 4.1.
@@ 3.4.2 Reformulation of GPD Parameter Estimators
To employ differentiable rank $\textbf{{r}}=r_{\varepsilon}(\textbf{{w}})$ for Pareto smoothing, since it represents ranks as continuous values, we need to modify the GPD parameter estimators, i.e., $\hat{\mu}$, $\hat{\sigma}$, and $\hat{\xi}$.
@@ 3.4.3 Overall Algorithm
Compared with the DRCFR method [Hassanpour and Greiner, 2020], our method requires additional time to compute Pareto-smoothed weights (lines 12-16 in Algorithm 1). In particular, computing differentiable rank (line 12) requires time complexity $\mathcal{O}(B\mathrm{log}B)$ for mini-batch size $B$, which is needed to evaluate the objective function in (3) and its gradient for each iteration in the training phase.
Remark: Strictly speaking, the choice of activation functions in propensity score $\pi$ and feature representations $\Gamma$ and $\Delta$ is critical for satisfying the assumption of Pareto smoothing that the distribution of the importance sampling weight is absolutely continuous, which is necessary to prove the asymptotic consistency (Theorem 1 of Vehtari et al. [2024]). This assumption holds if each activation is differentiable almost everywhere (i.e., differentiable except on a set of measure zero). However, for instance, using the rectified linear unit (ReLU) in propensity score model $\pi$ makes the distribution of IPW weight discontinuous, thus violating the assumption of Pareto smoothing. Even with almost everywhere differentiable activation functions, due to the lack of learning theory on neural network models, deriving the asymptotic consistency of our CATE estimator is extremely challenging and is left as our future work.
@@ B.1 Settings of Baselines
We evaluated the performance of the neural network methods (TARNet and GANITE), with the existing implementations [Curth and van der Schaar, 2021, Yoon et al., 2018].  [^4] https://github.com/AliciaCurth/CATENets ,  [^5] https://github.com/jsyoon0823/GANITE By contrast, we implemented the DRCFR method using PyTorch [Paszke et al., 2019].